# Cayley Graph Propagation

### v1

@@ 5 Experimentation
In this section, we empirically validate the efficacy of using the complete Cayley graph structure on a range of graph classification tasks. In particular, we compare the CGP model against approaches that do not incur a computational complexity of having to examine the input graph’s structure: master node [5], fully adjacent layer (FA) [6] and EGP [13]. In addition to these comparisons, we further examine the scalability and performance of CGP against state-of-the-art graph rewiring techniques [40, 8, 9, 41, 11], leveraging the TUDataset [42] and LRGB [43]. Due to space limitations, we defer the results of the LRGB to Appendix (Section B), as well as analysing over-smoothing by measuring the Dirichlet energy of a graph to Appendix (Section E).
@@
We evaluate on the Open Graph Benchmark (OGB) [44] and TUDataset [42]. In our experiments, we prioritise a fair comparison with the only modification to the model being the use of the complete Cayley graph structure and the initialisation of the virtual nodes. Furthermore, we show that our proposed method is MPNN invariant by setting the underlying model to GCN [45] and GIN [39]. The chosen hyperparameters are in line with established foundations [44, 9] for each dataset respectively. Notably, the EGP paper limits its empirical analysis to only the graph classification tasks from OGB with GIN as the backbone GNN. Refer to Appendix (Section A) for more detail on our experimental setting and hyperparameters used for the OGB and TUDataset.
@@
For real-world comparison and to extend the foundations of EGP, we first provide results on graph classification datasets from the OGB [44] against techniques that do not require dedicated preprocessing. ogbg-molhiv is among the largest molecule property prediction datasets within the scope of the MoleculeNet benchmark [46], thus providing emulation for real-world analysis. ogbg-ppa focuses on classifying species based on their taxa, using their protein-protein association networks [47]. Our model takes inspiration from the open-source implementation of OGB and hyperparameters as given by [44], including fixing the number of layers to 5, a hidden dimension of 300, a dropout of 50% and with the only modification being a batch size of 64. We report across 10 seeds and 5 seeds for ogbg-molhiv and ogbg-ppa respectively. Our results in Table 1 show that overall the CGP model outperforms the other approaches that do not require dedicated preprocessing; exemplified in the results for ogbg-molhiv. Moreover, CGP consistently outperforms the base GCN and GIN, which is not the case for the other baseline models that do not require dedicated preprocessing.
TUDataset
We extend our graph classification analysis by evaluating on REDDIT-BINARY, IMDB-BINARY, MUTAG, ENZYMES, PROTEINS and COLLAB from the TUDataset [42]. Significantly, these datasets were chosen under the claim of Karhadkar et al. [9] that the topology of the graphs in relation to the tasks require long-range interactions. The results reported in Table 2 are a continuation of our evaluation of CGP against the approaches that do not require dedicated preprocessing approaches as reported in Table 1. In contrast, the results reported in Table 3 are against the renowned graph rewiring techniques [40, 8, 9, 41, 11]. For both Table 2 and 3 we report the results averaged across 20 random seeds.
The hyperparameters used in Table 2 are similar to those prescribed by Karhadkar et al. [9], however, to further pinpoint with certainty that the performance gain can be credited to the utilisation of the complete Cayley graph structure we opt for those more akin to to our OGB experimentation. We set the number of layers to 5, the hidden dimension to 128 and use a dropout of 10%. The results in Table 2 underscore the effectiveness of our proposed model. For both GCN and GIN, CGP is consistently first or second across all datasets (except for one case), achieving the highest classification accuracy. This is particularly notable when the sparsity of the Cayley graphs is considered in relation to certain datasets, such as IMDB-BINARY and COLLAB. The sparse nature of the Cayley graph means that unlike many graph rewiring techniques edges may be removed; refer to Table 5 in Appendix (Section A) for the dataset statistics.
Table 2: Results of CGP compared against the baselines that do not require dedicated preprocessing for GCN and GIN on the TUDataset. OOM signifies out-of-memory on a NVIDIA RTX 4090. The colours highlight First, Second and Third.
@@
GCN ${\color[rgb]{1,0,0}90.000}\pm 2.121$ $63.200\pm 5.036$ $77.500\pm 10.186$ ${\color[rgb]{1,.5,0}39.083}\pm 6.465$ $71.696\pm 5.432$ $71.370\pm 2.394$
+ Master Node ${\color[rgb]{0,0,1}85.425}\pm 2.830$ ${\color[rgb]{0,0,1}64.600}\pm 5.342$ ${\color[rgb]{0,0,1}86.000}\pm 7.842$ $36.583\pm 5.904$ $72.277\pm 3.927$ ${\color[rgb]{1,0,0}73.590}\pm 1.694$
+ FA OOM $62.500\pm 5.296$ ${\color[rgb]{1,.5,0}84.500}\pm 7.890$ ${\color[rgb]{0,0,1}41.917}\pm 7.116$ ${\color[rgb]{0,0,1}73.259}\pm 4.397$ ${\color[rgb]{1,.5,0}71.980}\pm 2.413$
+ EGP ${\color[rgb]{1,.5,0}82.075}\pm 3.568$ ${\color[rgb]{1,.5,0}64.400}\pm 3.980$ $80.250\pm 7.980$ $38.917\pm 7.017$ ${\color[rgb]{1,.5,0}72.634}\pm 2.545$ $70.090\pm 2.225$
+ CGP $81.900\pm 2.691$ ${\color[rgb]{1,0,0}65.300}\pm 4.540$ ${\color[rgb]{1,0,0}89.250}\pm 8.983$ ${\color[rgb]{1,0,0}49.000}\pm 6.200$ ${\color[rgb]{1,0,0}74.330}\pm 3.992$ ${\color[rgb]{0,0,1}72.430}\pm 1.930$
GIN $89.500\pm 3.066$ ${\color[rgb]{1,.5,0}71.800}\pm 0.817$ $87.250\pm 9.284$ $44.583\pm 7.320$ $69.911\pm 4.792$ $71.410\pm 2.470$
+ Master Node ${\color[rgb]{1,0,0}91.600}\pm 1.868$ $70.400\pm 3.611$ ${\color[rgb]{1,.5,0}87.250}\pm 7.822$ $46.083\pm 6.396$ ${\color[rgb]{1,.5,0}73.170}\pm 3.793$ ${\color[rgb]{1,.5,0}72.920}\pm 2.394$
+ FA OOM ${\color[rgb]{1,0,0}73.150}\pm 3.554$ $85.500\pm 10.235$ ${\color[rgb]{1,0,0}55.083}\pm 5.974$ ${\color[rgb]{0,0,1}73.304}\pm 3.000$ ${\color[rgb]{0,0,1}73.310}\pm 1.720$
+ EGP ${\color[rgb]{1,.5,0}90.125}\pm 1.870$ $69.850\pm 4.028$ ${\color[rgb]{0,0,1}88.500}\pm 7.599$ ${\color[rgb]{1,.5,0}51.333}\pm 6.227$ $72.545\pm 4.319$ $72.820\pm 2.819$
+ CGP ${\color[rgb]{0,0,1}90.175}\pm 1.705$ ${\color[rgb]{0,0,1}72.200}\pm 2.804$ ${\color[rgb]{1,0,0}91.000}\pm 9.301$ ${\color[rgb]{0,0,1}52.667}\pm 6.442$ ${\color[rgb]{1,0,0}73.438}\pm 3.013$ ${\color[rgb]{1,0,0}74.420}\pm 2.202$
Table 3: Results of CGP compared against EGP and the graph rewiring techniques for GCN and GIN on the TUDataset. The experimental setup uses the setting as in Karhadkar et al. [9], and hyperparameters for each graph rewiring approach from [40, 8, 9, 41, 11]. The colours highlight First, Second and Third. OOT indicates out-of-time for the prepreprocessing rewiring time.
Model REDDIT-BINARY IMDB-BINARY MUTAG ENZYMES PROTEINS COLLAB
GCN $77.735\pm 1.586$ ${\color[rgb]{0,0,1}60.500}\pm 2.729$ $74.750\pm 4.030$ ${\color[rgb]{1,.5,0}29.083}\pm 2.363$ $66.652\pm 1.933$ ${\color[rgb]{1,.5,0}70.490}\pm 1.628$
@@
+ SDRF ${\color[rgb]{0,0,1}77.975}\pm 1.479$ $59.000\pm 2.254$ $74.000\pm 3.462$ $26.667\pm 2.000$ $67.277\pm 2.170$ ${\color[rgb]{0,0,1}71.330}\pm 0.807$
+ FoSR ${\color[rgb]{1,.5,0}77.750}\pm 1.385$ ${\color[rgb]{1,.5,0}59.750}\pm 2.357$ $75.250\pm 5.722$ $24.167\pm 3.005$ $70.848\pm 1.618$ $67.220\pm 1.367$
@@
+ GTR ${\color[rgb]{1,0,0}79.025}\pm 1.248$ ${\color[rgb]{1,0,0}60.700}\pm 2.079$ ${\color[rgb]{1,.5,0}76.500}\pm 4.189$ $25.333\pm 2.931$ ${\color[rgb]{1,.5,0}72.991}\pm 1.956$ ${\color[rgb]{1,0,0}72.600}\pm 1.025$
@@
GIN ${\color[rgb]{1,.5,0}84.600}\pm 1.454$ ${\color[rgb]{0,0,1}71.250}\pm 1.509$ $80.500\pm 5.143$ $35.667\pm 2.803$ $70.312\pm 1.749$ $71.490\pm 0.746$
+ DIGL $84.575\pm 1.265$ $52.650\pm 2.150$ $78.500\pm 4.189$ ${\color[rgb]{0,0,1}41.500}\pm 3.063$ ${\color[rgb]{0,0,1}72.321}\pm 1.440$ $57.620\pm 1.010$
+ SDRF $84.550\pm 1.396$ $69.550\pm 2.381$ ${\color[rgb]{1,.5,0}80.500}\pm 4.177$ $37.167\pm 2.709$ $69.509\pm 1.709$ ${\color[rgb]{0,0,1}72.958}\pm 0.419$
+ FoSR ${\color[rgb]{1,0,0}85.750}\pm 1.099$ $69.250\pm 1.810$ $80.500\pm 4.738$ $28.083\pm 2.301$ $71.518\pm 1.767$ $71.720\pm 0.892$
@@
+ GTR ${\color[rgb]{0,0,1}85.474}\pm 0.826$ $69.550\pm 1.473$ $79.000\pm 3.847$ $31.750\pm 2.466$ ${\color[rgb]{1,.5,0}72.054}\pm 1.510$ $71.849\pm 0.710$
+ EGP $77.875\pm 1.563$ $68.250\pm 1.121$ ${\color[rgb]{0,0,1}81.500}\pm 4.696$ ${\color[rgb]{1,.5,0}40.667}\pm 3.095$ $70.848\pm 1.568$ ${\color[rgb]{1,.5,0}72.330}\pm 0.954$
@@
In Table 3 we compare CGP against the baselines, including the state-of-the-art graph rewiring techniques: DIGL [40], SDRF [8], FoSR [9], BORF [41] and GTR [11]. However, in this instance we use the experimental setting and hyperparameters set by Karhadkar et al. [9], which are commonly used across the aforementioned graph rewiring techniques. By applying these hyperparameters, we not only ensure a fair comparison, but showcase that CGP continues to be competitive even under their conditions. The hyperparameters include a hidden dimension of 64, the number of layers set to 4 and a dropout of 50%. Finally, Karhadkar et al. [9] reports the results with a 95% confidence interval. Thus, we respect this for decision Table 3. Notably, we report OOT to indicate out-of-time for the preprocessing rewiring procedure for BORF on the REDDIT-BINARY and COLLAB datasets. This is in accordance with Nguyen et al. [41] who do not report results for these two datasets and corresponding hyperparameters. In addition, a time-out is reported in [48] for the aforementioned datasets, whilst in [49] they report out-of-memory for COLLAB. In Appendix (Section C) we conduct a scalability analysis in which we provide the graph rewiring time for each of our baseline models, as well as extending this beyond the real-world datasets through a synthetic benchmark.
@@
The results in Table 3 underscore the effectiveness of CGP in comparison with the graph rewiring techniques. In particular, in the case of GIN, our CGP model obtains the highest accuracy for all datasets except for REDDIT-BINARY. The overall performance of CGP when applied to GCN is not as competitive as those of GIN, however our results are still comparable with the other graph rewiring techniques. CGP obtains the highest accuracy in two of the datasets, only placing second to GTR. Collectively, our results are consistent with our previously discussed limitations regarding the sparsity of CGP when applied to datasets with a significantly higher edge-to-node ratio. However, the results of CGP for GIN recover this loss in performance for IMDB-BINARY and COLLAB, emphasising the work of You et al. [50] in which the design space of GNNs can greatly impact a model’s results. This is supplemented by the results reported in Table 2 in which IMDB-BINARY for GCN obtains the highest classification accuracy. Finally, of significance is the parity of the hyperparameter’s number of layers; Table 2 uses 5 layers, whereas Table 3 uses 4 layers. This demonstrates the performance of CGP is irrespective of the final layer being the input or Cayley graph.
@@
In the following, we answer the question: ‘is the complete Cayley graph structure a suitable alternative to a fully adjacent layer [6]?’. In Table 4 we show that it is a promising avenue as the results for CGP and FA are similar for both GCN and GIN. However, the sparsity of CGP is highlighted in the results of REDDIT-BINARY, whereby FA runs out-of-memory (OOM). Accordingly, these results demonstrate the advantages of CGP, due to it being far more scalable. Notably, the results reported in Table 4 use the same experimental setting and hyperparameters as Table 2.
Next, we investigate in accordance with the recent work of Bechler-Speicher et al. [51]; we examine if we can ignore the input graph entirely and solely propagate over the Cayley graph structure. Our results in Table 4 conclude that the inductive bias endowed from the input graph is required. The baseline procedure of CGP (interweaving a Cayley graph with an input graph), and using a Cayley graph for the last layer only, outperforms using a Cayley graph solely for each layer. One explanation is that the interweaving schema of CGP aligns with the principles of JK networks [52], facilitating improved structure-aware representations by varying the neighbourhood ranges. However, the tone set by Bechler-Speicher et al. [51] is still a promising line of research, as the results of PROTEINS indicates that the optimal graph-structure used is still task dependent.
@@
GCN ${\color[rgb]{1,0,0}90.000}\pm 2.121$ ${\color[rgb]{0,0,1}63.200}\pm 5.036$ $77.500\pm 10.186$ $39.083\pm 6.465$ $71.696\pm 5.432$ ${\color[rgb]{1,.5,0}71.370}\pm 2.394$
+ FA OOM ${\color[rgb]{1,.5,0}62.500}\pm 5.296$ ${\color[rgb]{1,.5,0}84.500}\pm 7.890$ ${\color[rgb]{1,.5,0}41.917}\pm 7.116$ ${\color[rgb]{1,.5,0}73.259}\pm 4.397$ ${\color[rgb]{0,0,1}71.980}\pm 2.413$
+ FA OOM $48.800\pm 6.547$ $81.500\pm 9.631$ $29.667\pm 4.989$ $71.786\pm 2.452$ $42.940\pm 14.368$
+ CGP ${\color[rgb]{1,.5,0}81.900}\pm 2.691$ ${\color[rgb]{1,0,0}65.300}\pm 4.540$ ${\color[rgb]{1,0,0}89.250}\pm 8.983$ ${\color[rgb]{1,0,0}49.000}\pm 6.200$ ${\color[rgb]{1,0,0}74.330}\pm 3.992$ ${\color[rgb]{1,0,0}72.430}\pm 1.930$
+ CGP ${\color[rgb]{0,0,1}89.600}\pm 1.736$ $61.850\pm 5.850$ $78.750\pm 9.601$ ${\color[rgb]{0,0,1}42.000}\pm 4.522$ $71.250\pm 3.989$ $71.360\pm 1.663$
+ CGP $70.250\pm 3.673$ $54.550\pm 4.295$ ${\color[rgb]{0,0,1}88.500}\pm 7.921$ $35.750\pm 5.664$ ${\color[rgb]{0,0,1}73.259}\pm 3.346$ $56.170\pm 2.412$
GIN ${\color[rgb]{1,.5,0}89.500}\pm 3.066$ $71.800\pm 0.817$ ${\color[rgb]{1,.5,0}87.250}\pm 9.284$ $44.583\pm 7.320$ $69.911\pm 4.792$ $71.410\pm 2.470$
+ FA OOM ${\color[rgb]{1,0,0}73.150}\pm 3.554$ $85.500\pm 10.235$ ${\color[rgb]{1,0,0}55.083}\pm 5.974$ ${\color[rgb]{1,.5,0}73.304}\pm 3.000$ ${\color[rgb]{1,.5,0}73.310}\pm 1.720$
@@
+ CGP ${\color[rgb]{0,0,1}90.175}\pm 1.705$ ${\color[rgb]{1,.5,0}72.200}\pm 2.804$ ${\color[rgb]{1,0,0}91.000}\pm 9.301$ ${\color[rgb]{0,0,1}52.667}\pm 6.442$ ${\color[rgb]{0,0,1}73.438}\pm 3.013$ ${\color[rgb]{1,0,0}74.420}\pm 2.202$
+ CGP ${\color[rgb]{1,0,0}90.200}\pm 1.749$ ${\color[rgb]{0,0,1}72.900}\pm 4.381$ ${\color[rgb]{0,0,1}90.000}\pm 5.701$ ${\color[rgb]{1,.5,0}51.417}\pm 6.714$ $71.741\pm 4.667$ ${\color[rgb]{0,0,1}73.760}\pm 1.664$
+ CGP $75.775\pm 3.378$ $53.250\pm 4.276$ $85.500\pm 7.730$ $37.167\pm 6.261$ ${\color[rgb]{1,0,0}74.330}\pm 4.177$ $57.820\pm 2.172$

### v2

@@ 5 Experimentation
In this section, we empirically validate the efficacy of using the complete Cayley graph structure on a range of graph classification tasks. In particular, we first motivate CGP by extending the results of Deac et al. [14], comparing CGP against approaches that do not incur a computational complexity of having to examine the input graph’s structure: master node [5], fully adjacent last layer (FA) [6] and EGP [14]. We then extend our empirical evaluation by comparing the performance of CGP against the state-of-the-art approaches that require dedicated preprocessing [33, 9, 10, 34, 12, 35] on the TUDataset [36] and LRGB [37]. Due to space limitations, we defer the results of the LRGB to Appendix (Section D).
Beyond comparing the performance of CGP against the baselines, we evaluate the practicality of CGP by providing a runtime analysis against the aforementioned approaches that require dedicated processing. In Appendix (Section E), we extend this through a scalability analysis. Furthermore, we provide an ablation study to investigate whether the complete Cayley graph structure is a suitable alternative to a fully-connected graph.
@@
We evaluate on the Open Graph Benchmark (OGB) [38] and TUDataset [36]. In our experiments, we prioritise a fair comparison, following the layer schema for each approach. For EGP and CGP, this includes the interweaving schema as depicted in Section 4. For FA [6], this consists of rewiring the last layer. All the other approaches only propagate over the base (or rewired) graph structure. Furthermore, we show that our proposed method is MPNN invariant by setting the underlying model to GCN [39] and GIN [32]. The chosen hyperparameters are in line with established foundations [38, 10] for each dataset respectively. Notably, EGP [14] limits its empirical analysis to only the graph classification tasks from OGB with GIN as the backbone GNN. Refer to Appendix (Section C) for more details on our experimental setting and hyperparameters used for the OGB and TUDataset.
@@
For a real-world comparison and to extend the foundations of EGP, we first provide results on two graph classification tasks, ogbg-molhiv and ogbg-ppa, from the OGB [38]. We compare CGP against techniques that do not require dedicated preprocessing. ogbg-molhiv is among the largest molecule property prediction datasets within the scope of the MoleculeNet benchmark [40], thus providing emulation for real-world analysis. ogbg-ppa focuses on classifying species based on their taxa, using their protein-protein association networks [41]. Our model takes inspiration from the open-source implementation of OGB and hyperparameters as given by [38], including fixing the number of layers to 5, a hidden dimension of 300, a dropout of 50% and with the only modification being a batch size of 64. We report across 10 seeds and 5 seeds for ogbg-molhiv and ogbg-ppa respectively. Our results in Table 1 show that overall the CGP model outperforms the other approaches that do not require dedicated preprocessing; exemplified in the results for ogbg-molhiv. Moreover, CGP consistently outperforms the base GCN and GIN, which is not the case for the other baseline models.
Table 2: Results of CGP compared against EGP, FA and the approaches that require dedicated preprocessing for GCN and GIN on the TUDataset. The experimental setup uses the setting as in Karhadkar et al. [10], and hyperparameters for each baseline from [33, 9, 10, 34, 12, 35]. The colours highlight First, Second and Third. OOT indicates out-of-time for the dedicated prepreprocessing time.
@@
GCN $77.735\pm 1.586$ ${\color[rgb]{1,.5,0}60.500}\pm 2.729$ $74.750\pm 4.030$ $29.083\pm 2.363$ $66.652\pm 1.933$ $70.490\pm 1.628$
+ FA OOM $48.950\pm 1.652$ $70.250\pm 4.608$ $28.667\pm 3.693$ $71.071\pm 1.506$ ${\color[rgb]{1,.5,0}72.039}\pm 0.771$
@@
+ SDRF ${\color[rgb]{1,.5,0}77.975}\pm 1.479$ $59.000\pm 2.254$ $74.000\pm 3.462$ $26.667\pm 2.000$ $67.277\pm 2.170$ $71.330\pm 0.807$
+ FoSR $77.750\pm 1.385$ $59.750\pm 2.357$ $75.250\pm 5.722$ $24.167\pm 3.005$ $70.848\pm 1.618$ $67.220\pm 1.367$
@@
+ GTR ${\color[rgb]{0,0,1}79.025}\pm 1.248$ ${\color[rgb]{0,0,1}60.700}\pm 2.079$ $76.500\pm 4.189$ $25.333\pm 2.931$ ${\color[rgb]{1,.5,0}72.991}\pm 1.956$ ${\color[rgb]{0,0,1}72.600}\pm 1.025$
+ PANDA ${\color[rgb]{1,0,0}87.275}\pm 1.033$ ${\color[rgb]{1,0,0}68.350}\pm 2.346$ ${\color[rgb]{1,.5,0}76.750}\pm 5.531$ ${\color[rgb]{1,.5,0}30.667}\pm 2.019$ $70.134\pm 1.518$ ${\color[rgb]{1,0,0}73.850}\pm 0.695$
@@
GIN $84.600\pm 1.454$ ${\color[rgb]{0,0,1}71.250}\pm 1.509$ $80.500\pm 5.143$ $35.667\pm 2.803$ $70.312\pm 1.749$ $71.490\pm 0.746$
+ FA OOM $69.900\pm 2.332$ $80.250\pm 5.314$ ${\color[rgb]{0,0,1}47.833}\pm 2.529$ ${\color[rgb]{0,0,1}72.902}\pm 1.419$ $72.740\pm 0.786$
+ DIGL $84.575\pm 1.265$ $52.650\pm 2.150$ $78.500\pm 4.189$ $41.500\pm 3.063$ ${\color[rgb]{1,.5,0}72.321}\pm 1.440$ $57.620\pm 1.010$
+ SDRF $84.550\pm 1.396$ $69.550\pm 2.381$ $80.500\pm 4.177$ $37.167\pm 2.709$ $69.509\pm 1.709$ ${\color[rgb]{1,.5,0}72.958}\pm 0.419$
+ FoSR ${\color[rgb]{0,0,1}85.750}\pm 1.099$ $69.250\pm 1.810$ $80.500\pm 4.738$ $28.083\pm 2.301$ $71.518\pm 1.767$ $71.720\pm 0.892$
@@
+ GTR ${\color[rgb]{1,.5,0}85.474}\pm 0.826$ $69.550\pm 1.473$ $79.000\pm 3.847$ $31.750\pm 2.466$ $72.054\pm 1.510$ $71.849\pm 0.710$
+ PANDA ${\color[rgb]{1,0,0}90.325}\pm 0.867$ $68.350\pm 2.346$ ${\color[rgb]{0,0,1}83.250}\pm 3.262$ ${\color[rgb]{1,.5,0}42.167}\pm 2.286$ $72.321\pm 1.786$ ${\color[rgb]{0,0,1}73.320}\pm 0.814$
+ EGP $77.875\pm 1.563$ $68.250\pm 1.121$ ${\color[rgb]{1,.5,0}81.500}\pm 4.696$ $40.667\pm 3.095$ $70.848\pm 1.568$ $72.330\pm 0.954$
@@
TUDataset
We extend our graph classification task analysis by evaluating on REDDIT-BINARY, IMDB-BINARY, MUTAG, ENZYMES, PROTEINS and COLLAB from the TUDataset [36]. Significantly, these datasets were chosen under the claim of Karhadkar et al. [10] that the topology of the graphs in relation to the tasks require long-range interactions. Accordingly, the TUDataset has become the cornerstone collection of benchmark datasets for the graph-rewiring approaches investigating over-squashing. Thus, we compare CGP against EGP [14] and FA [29], as well as the following state-of-the-art graph rewiring techniques that require dedicated preprocessing: DIGL [33], SDRF [9], FoSR [10], BORF [34] and GTR [12]. Moreover, we include PANDA [35] as a unique approach that dynamically alters the width [7] of the model to alleviate over-squashing as opposed to rewiring the input graph structure.
The results in Table 2 prioritise a fair comparison to further pinpoint with certainty that the performance gain can be credited to the utilisation of the complete Cayley graph structure. Therefore, in Table 2 we use the hyperparameters and experimental setting as prescribed by Karhadkar et al. [10]. The hyperparameters include a hidden dimension of 64, the number of layers set to 4 and a dropout of 50%. In line with our OGB experimentation for all models, we also use Batch Norm [42]. Unlike the baseline model, FA, EGP and CGP, the dedicated preprocessing approaches have an optimisation target and thus feature additional approach specific hyperparameters. For each dataset, we use the optimised hyperparameter setting as stated in each paper respectively. More details are found in Appendix (Section C). The reported results are averaged across 20 random seeds and Karhadkar et al. [10] reports the results with a 95% confidence interval, thus we respect this for the TUDataset. Notably, we report OOT to indicate out-of-time for the preprocessing rewiring procedure for BORF on the REDDIT-BINARY and COLLAB datasets. This is in accordance with Nguyen et al. [34], which do not report results for these two datasets and corresponding hyperparameters. In addition, a time-out is reported in [35] for the aforementioned datasets, whilst in [43] they report out-of-memory for COLLAB.
@@
The results in Table 2 underscore the effectiveness of CGP in comparison with state-of-the-art baselines. In particular, in the case of GIN, our CGP model obtains the highest accuracy for all datasets except for REDDIT-BINARY. The overall performance of CGP when applied to GCN is not as competitive as those of GIN, however our results are still comparable with the other baselines. This is particularly notable when the sparsity of the Cayley graphs is considered in relation to certain datasets, such as IMDB-BINARY and COLLAB. The sparse nature of the Cayley graph means that unlike many graph rewiring techniques edges may be removed; refer to Appendix (Section C) for the dataset statistics. However, the results of CGP for GIN recover this loss in performance for IMDB-BINARY and COLLAB, emphasising the work of You et al. [44] in which the design space of GNNs can greatly impact the results of a model. Finally, of significance is the parity of the hyperparameter’s number of layers; Table 1 uses 5 layers, whereas Table 2 uses 4 layers. This demonstrates the performance of CGP is irrespective of the final layer being the input or Cayley graph.
Table 3: CGP training, evaluation time (seconds per epoch), and memory consumption statistics in comparison to baseline models on REDDIT-BINARY and COLLAB from the TUDataset [36]. OOM signifies out-of-memory on a NVIDIA RTX 4090.
Model REDDIT-BINARY COLLAB
Train Time Eval. Time Mem. (GB) Train Time Eval. Time Mem. (GB)
GIN $0.1049\pm 0.0237$ $0.0741\pm 0.0032$ $922$ $0.2787\pm 0.0345$ $0.2364\pm 0.0094$ $1722$
+ FA OOM OOM OOM $0.4625\pm 0.0507$ $0.4488\pm 0.0404$ $4746$
+ FoSR $0.1117\pm 0.0268$ $0.0841\pm 0.0164$ $906$ $0.3129\pm 0.0386$ $0.2619\pm 0.0257$ $3320$
+ PANDA $0.7902\pm 0.0597$ $0.7489\pm 0.0439$ $1316$ $2.1152\pm 0.0964$ $1.9347\pm 0.0924$ $4406$
+ EGP $0.1215\pm 0.0257$ $0.0952\pm 0.0128$ $976$ $0.3096\pm 0.0372$ $0.2598\pm 0.0164$ $1696$
+ CGP $0.1326\pm 0.0296$ $0.1147\pm 0.0135$ $1128$ $0.3191\pm 0.0321$ $0.2785\pm 0.0160$ $2418$
Scalability
In the following, we investigate whether the additional virtual nodes as leveraged in CGP will introduce an increased computational complexity, impacting the training time of the model. Accordingly, we compare CGP directly against EGP [14], FA [6] and the graph-rewiring approach FoSR [10]. Moreover, we include PANDA to provide extra detail on the results found in Table 2. Even though PANDA obtains state-of-the-art performance, as stated by Choi et al. [35] this approach impacts the runtime. We choose REDDIT-BINARY and COLLAB from the TUDataset because they have the largest average graph sizes among the collection of datasets. In Table 3 we report the average seconds per epoch for the training and evaluation time, as well as the memory consumption statistics, using GIN as the underlying model.
The results highlight that additional virtual nodes in CGP have a negligible impact on the training and evaluation time. Even though many virtual nodes may be added, they are sparsely connected. This is depicted in Figure 1, where the 4-regular structure of the Cayley graph results in virtual nodes being sparsely connected compared to other virtual node approaches, such as a master node [5]. The base model, FoSR [10], EGP [14] and CGP are shown to be akin with each other. CGP is shown to increase the memory consumption, however the sparsely connected virtual nodes have a minimal impact when compared with FA [29], FoSR [10] and PANDA [35]. The dataset statistics in Appendix (Section C) provide an explanation, such as REDDIT-BINARY being several times larger than all other datasets found in the TUDataset [36]. Overall, the results in Table 3 highlight the strengths of CGP against the approaches that require dedicated preprocessing, as well as the dense fully adjacent layer [29]. In Appendix (Section E), we extend our scalability analysis of CGP. This includes conducting a scalability analysis to compare the dedicated preprocessing time of the techniques reported in Table 2 on the real-world datasets from the TUDataset, as well as further extending this evaluation through a synthetic benchmark as used by Karhadkar et al. [10].
@@
In the following, we answer the question: ‘is the complete Cayley graph structure a suitable alternative to a fully adjacent layer [6]?’. In Table 4, we show that it is a promising avenue as the results for CGP and FA are similar for both GCN and GIN. However, the sparsity of CGP is highlighted in the results of REDDIT-BINARY, whereby FA runs out-of-memory (OOM). Accordingly, these results demonstrate the advantages of CGP, due to it being far more scalable. Notably, the results reported in Table 4 use the same experimental setting and hyperparameters as Table 2.
Next, we investigate in accordance with the recent work of Bechler-Speicher et al. [45]; we examine if we can ignore the input graph entirely and solely propagate over the Cayley graph structure. Our results in Table 4 suggest that the inductive bias endowed from the input graph still is required. The baseline procedure of CGP (interweaving a Cayley graph with an input graph), and using a Cayley graph for the last layer only, outperforms using a Cayley graph solely for each layer. One explanation is that the interweaving schema of CGP aligns with the principles of JK networks [46], facilitating improved structure-aware representations by varying the neighbourhood ranges. However, the tone set by Bechler-Speicher et al. [45] is still a promising line of research, as the results indicate that the optimal graph-structure used is still task dependent.
@@
GCN ${\color[rgb]{1,0,0}77.735}\pm 1.586$ ${\color[rgb]{1,0,0}60.500}\pm 2.729$ $74.750\pm 4.030$ ${\color[rgb]{1,.5,0}29.083}\pm 2.363$ $66.652\pm 1.933$ ${\color[rgb]{0,0,1}70.490}\pm 1.628$
+ FA OOM $48.950\pm 1.652$ $70.250\pm 4.608$ $28.667\pm 3.693$ $71.071\pm 1.506$ ${\color[rgb]{1,0,0}72.039}\pm 0.771$
+ FA OOM $49.700\pm 1.871$ $75.250\pm 4.554$ $27.167\pm 2.770$ ${\color[rgb]{0,0,1}71.384}\pm 1.380$ $54.990\pm 0.699$
+ CGP ${\color[rgb]{0,0,1}67.050}\pm 1.483$ ${\color[rgb]{0,0,1}56.200}\pm 1.825$ ${\color[rgb]{1,0,0}83.750}\pm 3.597$ ${\color[rgb]{1,0,0}31.000}\pm 2.397$ ${\color[rgb]{1,0,0}73.036}\pm 1.291$ ${\color[rgb]{1,.5,0}69.630}\pm 0.730$
+ CGP ${\color[rgb]{1,.5,0}56.025}\pm 2.185$ $51.200\pm 1.729$ ${\color[rgb]{0,0,1}77.750}\pm 3.706$ ${\color[rgb]{0,0,1}30.167}\pm 3.196$ $66.875\pm 2.441$ $59.050\pm 1.077$
+ CGP $54.949\pm 1.531$ ${\color[rgb]{1,.5,0}52.350}\pm 2.242$ ${\color[rgb]{1,.5,0}76.000}\pm 4.219$ $28.918\pm 3.191$ ${\color[rgb]{1,.5,0}68.124}\pm 3.346$ $58.450\pm 1.321$
GIN ${\color[rgb]{1,0,0}84.600}\pm 1.454$ $71.250\pm 1.509$ $80.500\pm 5.143$ $35.667\pm 2.803$ $70.312\pm 1.749$ $71.490\pm 0.746$
+ FA OOM $69.900\pm 2.332$ $80.250\pm 5.314$ $47.833\pm 2.529$ ${\color[rgb]{0,0,1}72.902}\pm 1.419$ ${\color[rgb]{1,.5,0}72.740}\pm 0.786$
@@
+ CGP $78.225\pm 1.268$ ${\color[rgb]{1,0,0}71.650}\pm 1.532$ ${\color[rgb]{1,.5,0}85.250}\pm 3.200$ ${\color[rgb]{0,0,1}50.083}\pm 2.242$ ${\color[rgb]{1,0,0}73.080}\pm 1.396$ ${\color[rgb]{1,0,0}73.350}\pm 0.788$
+ CGP ${\color[rgb]{0,0,1}82.700}\pm 2.474$ ${\color[rgb]{0,0,1}71.500}\pm 1.646$ ${\color[rgb]{1,0,0}88.500}\pm 4.955$ ${\color[rgb]{1,0,0}52.000}\pm 3.355$ ${\color[rgb]{1,.5,0}71.383}\pm 2.472$ ${\color[rgb]{0,0,1}72.880}\pm 0.834$
+ CGP ${\color[rgb]{1,.5,0}80.925}\pm 1.976$ ${\color[rgb]{1,.5,0}71.400}\pm 2.002$ ${\color[rgb]{0,0,1}85.500}\pm 3.309$ ${\color[rgb]{1,.5,0}49.167}\pm 3.184$ $70.937\pm 1.788$ $72.340\pm 1.038$